Prev Next

SW&IT Meeting Minutes 2021-01-11

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20201228
LR20201228 - Bec, Matthieu 40.0
LR20201228 - Edmeades, David 40.0
LR20201228 - Sola, Igor 0.0
LR20201228 - Grenz, Paul 0.0
LR20201228 - Hooper, Stephen 0.0
LR20201228 - Bryant, Trey 0.0
LR20201228 - Zhang, Yang 0.0
LR20201228 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT SW / ATT PM - documents shipped
    • IT7917 and IT7923 (Vx socket error, LBTI AO)
    • script to automate weekly minutes
    • adjust TMS alarm handler, issue with powermeter turning off
    • new TCS/AO SW job desc.
  • This week
    • new SW student job desc.
    • SHARK-NIR shutdown and emergency
    • resume DMS websocket gateway
    • kickstart swgdev in various ways

YZ

  • Previous:
    • Vacation
  • Next:
    • Catch up with some recent emails and activities.
    • Attend TMS and AO-Cravat meetings and discussions.
    • Continue with OSCO web prototype.

DE

  • post-holiday catchup
  • Archive, various
    • Cristina assures me that she's working on pushing data out of /newdata. A check this morning bears this out, current usage 4%
  • Bucki accounts
  • Received a replacement HD
  • Cisco
Next:
  • Hearing issues re: wired registration in Steward, though my office workstation is accessable. Let us know if you have problems.
  • Synology disk 5 is complaining more, should take care of it next visit. It's not currently bad enough to schedule a visit just for that
  • Heads up next week we have work in 573. I don't anticipate any downtime.
  • Cisco

SH

  • JCG stuff
  • LBTI and sw0603 troubleshooting for sxvx
  • CPE to UITS
  • Microwave ticket from 12/25 just got closed

next:

  • work on mountain microwave (pretty much not for me)
  • e-mail PEPSI
  • Steward CR meeting
  • robs2

IS

Last week:
  • AO/ARGOS documentation
  • LUCI
    • IT #7493 Aladin slow subtraction
    • IT #8231- Timeout on GEIRS during readout
  • PIT
    • Working on a proposals table.

This week:
  • AO/ARGOS documentation
  • LUCI
    • IT #7493 Aladin slow subtraction
    • IT #8231- Timeout on GEIRS during readout
    • IT #8026- Re-measure LUCI1 AO patrol field and update OT
  • PIT
    • Populate table with real and mock up data

PSG

  • Last week
    • Testing White light source
  • This week
    • Further testing of swingarm
    • Determining why error occurs when motion is requested.

XZ

  • Last week
    • vacation
    • ATT AdSec and SW documents
  • This week
    • AdSec safety SW test
    • Work/update on few ITs
    • AO support
    • Blind deconvolution test
Topic revision: r11 - 11 Jan 2021, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback